# Phase diagram of the interacting Haldane model with spin-dependent sublattice potentials

### v1

@@ III.1 Results of the Exact Diagonalization
To detail the critical behaviors of the phase diagram, we select two balck dashed lines with $V=0.4$ and $\Delta=1.4$ in Fig. 1(a) to calculate the relevant properties. The first four lowest-lying energy levels $E_{\alpha}$ ($\alpha=0,1,2,3$ and $\alpha=0$ represents the ground state) are obtained by employing the Arnoldi Lehoucq et al. (1997) method, see Figs. 2(a) and 2(e). In Fig. 2(a) with $V=0.4$, the system crosses the $C=2$, $C=1$ and SDW phases as $\Delta$ increases from $0.6$ to $1.4$, and two level crossings between the ground state and one excited state can be observed at $\Delta\approx 0.9$ and $\Delta\approx 1.2$. In Fig. 2(e) with $\Delta=1.4$, the system crosses the SDW, $C=1$ and CDW phases as $V$ increases from $0.0$ to $2.0$, and two level crossings locates at $V\approx 0.75$ and $V\approx 1.25$. It may be more intuitive to exhibit the excitation gaps $\Delta_{\rm ex}^{1}=E_{1}-E_{0}$ and $\Delta_{\rm ex}^{2}=E_{2}-E_{0}$, see Figs. 2(b) and 2(f). We can find that the excitation gaps exhibit local minimum values at the critical points where level crossings occur. It should be pointed out that in Fig. 2(f), $\Delta_{\rm ex}^{1}$ becomes very small when $V>1.25$, which is due to the almost degenerate ground state and first excited state in CDW phase. Another feature to characterize the phase transitions is the change of the CDW or SDW structure factor ($S_{\text{CDW}}$ or $S_{\text{SDW}}$) in Figs. 2(c) and 2(g). Values of $S_{\text{CDW}}$ ($S_{\text{SDW}}$) in $C=1$ phase is larger than those in $C=2$ phase but smaller than those in CDW (SDW) phase, which indicating that $C=1$ is an intermediate state as a result of the interplay between topology, electronic correlations and lattice potentials. In addition, sharp peaks of the ground-state fidelity metrics $g$ in Figs. 2(d) and 2(h) can be also observed and used to characterize the critical points. Finally, notice that the left and right panels in Fig. 2 are corresponding to the results of $V=0.4$ and $\Delta=1.4$, respectively.
For $U=3.0$ in Fig. 1(d), similar calculations and discussions along the black dashed lines of $\Delta=0.4$ and $\Delta=0.6$ are detailed in Appendix A.
@@ III.2 Results of the Mean-Field Method
To contrast the exact results on small lattice above, we now report the outcomes of the MF method. By implementing a variational mean-field approximation, the lattice size can be enlarged significantly. We use the $180\times 180$ lattice to calculate the band gap and the order parameters ($\mathcal{O}_{\text{CDW}}$ and $\mathcal{O}_{\text{SDW}}$), while a $30\times 30$ lattice is adopted for the calculations of Chern number. Even though their lattice size are different, critical points identified by the Chern number and other properties are consistent, as will be discussed in Fig. 4. Phase diagrams with regard to the parameters $V$ and $\Delta$ are shown in Figs. 3(a), 3(b), 3(c) and 3(d) with $U=0.0$, $U=1.0$, $U=2.0$ and $U=3.0$, respectively. Compared to the phase diagrams of ED in Fig. 1, similar features can be observed: SDW and CDW dominate the system for large $\Delta$ and $V$, respectively, leaving the $C=2$ phase in small ($V$, $\Delta$) region; the $C=1$ phase are surrounded by the above three phases. Differently, the increasing interaction $U$ suppresses but does not eliminate the $C=1$ state.
@@
Figure 3: Mean-field phase diagram in the parametric space ($V$, $\Delta$) of the model (1) based on the results of Chern number $C$ and the order parameters ($\mathcal{O}_{\text{CDW}}$ and $\mathcal{O}_{\text{SDW}}$), with (a) $U=0.0$, (b) $U=1.0$, (c) $U=2.0$ and (d) $U=4.0$.
@@
Taking the $U=0.0$ case in Fig. 3(a) as an example, we show in Figs. 4(a) and 4(b) the CDW and SDW order parameters ($\mathcal{O}_{\text{CDW}}$ and $\mathcal{O}_{\text{SDW}}$), respectively, as a function of $V$ and $\Delta$. We find that $\mathcal{O}_{\text{CDW}}$ vanishes in SDW and $C=2$ phases, while the $C=1$ phase can be regarded as a zone of transition for $\mathcal{O}_{\text{CDW}}$ decreases from $2$ to 00. Notice that the phase boundary between SDW and CDW can only be identified from the order parameters, and other phase boundaries are characterized by the results of Chern number, see Fig. 4(d). Even though there are some defect points in Fig. 4(d) which may be due to the severe quantum fluctuation in the vicinity of the critical points, the phase boundaries can be readily identified. For $\mathcal{O}_{\text{SDW}}$ in Fig. 4(b), the $C=1$ phase is still a transition region with finite values but smaller that those in SDW phase. These features are similar to the structure factors of ED results shown in Figs. 2(c) and 2(g). To contrast the excitation gaps in ED results, we show the band gap defined as $\Delta(\textbf{k})$=min [$E_{2}(\textbf{k}^{\prime})-E_{1}(\textbf{k}^{\prime})$] in Fig. 4(c). The rapid drop or closing of the gap size can be observed at the critical values except for the phase boundary between SDW and CDW, where only discontinuity of gap size appears instead. This can be connected with the general picture that the change of a topological invariant is always accompanied by a single-particle gap closing. For the phase transition from SDW to CDW side, the Chern number does not change ($C=0$). That may be a reason to explain the discontinuity of gap size instead of gap closing, and remember that they are all gapped phases with different gap sizes.
@@
Figure 4: Contour plots of (a) the CDW order parameter, (b) the SDW order parameter, (c) the band gap and (d) the Chern number, as a function of $V$ and $\Delta$. The mean-field approach is adopted for the calculations of the model (1) with $U=0.0$.
@@ IV Summary and Discussion
To summarize, we studied the interacting spinful Haldane model at half-filling on the honeycomb lattice with spin-dependent sublattice potentials. By employing the exact-diagonalization (ED) and mean-field (MF) method, we obtained the similar ground-state phase diagrams with regard to the nearest-neighbor interaction $V\in[0,3]$ and sublattice potential difference $\Delta\in[0,3]$: the CDW and SDW phases are in the large $V$ and $\Delta$ region, respectively; the $C=2$ phase dominates the system when both $V$ and $\Delta$ are small enough; a intermediate phase with $C=1$ are surrounded by the above three phases. Except for the change of topological invariant, i.e., the Chern number, other features like the closure of excitation gap, change of structure factors and peak of the fidelity metric in ED results are also observed. As a contrast, the mean-field Chern number, band gap and local order parameters are also studied, which makes the existence of such $C=1$ phase more evident.
On the other hand, the interaction $U$ suppresses the $C=1$ phase but leaving an open question of whether it can exist with $U$ increased to $3.0$. In ED, $C=1$ can not be found when $U=3.0$ even though an intermediate phase can still be identified by results of the fidelity metric. In MF, $C=1$ phase preserves for $U=3.0$ even though its area is very small. Further studies on this interesting issue are thus expected.
@@ Acknowledgements.
C. S. acknowledges support from the National Natural Science Foundation of China (NSFC; Grant No. 12104229) and the Fundamental Research Funds for the Central Universities (Grant No. 30922010803). H.-G. L. acknowledges support from NSFC (Grants No. 11834005 and No. 12247101), and the National Key Research and Development Program of China (Grant No. 2022YFA1402704).
Appendix A details of some properties with $U=3$
@@
Figure 5: (a)(e) Four lowest-lying energy levels $E_{\alpha}$, (b)(f) the excitation gaps $\Delta_{\text{ex}}^{(\alpha)}$, (c)(g) the structure factors $S_{\text{SDW/CDW}}$, and (d)(h) the fidelity metric $g$ of the model (1) with $\Delta=0.4$ on the left panels and $\Delta=0.6$ on the right panels. The on-site interaction $U=3$ and the parameters are corresponding to the balck dashed lines in Fig. 1(d).
@@
In the case of $U=3.0$, we select $\Delta=0.4$ (left panel) and $\Delta=0.6$ (right panel) to present the relevant properties in Fig. 5. Remember that we do not observe the results of $C=1$ in Fig. 1(d) with $U=3.0$. The line of $\Delta=0.4$ ($\Delta=0.6$) crosses the $C=2$ phase (SDW phase), the intermediate phase and the CDW phase as $V$ increases from $1.0$ to $2.0$ [see black dashed lines in Fig. 1(d)]. From the ground-state fidelity metrics $g$ in Figs. 5(d) and 5(h), we can find the sharp peaks between the intermediate and the CDW phases. Accompanied with the shape peaks, level crossings between the ground state and second excited state [Figs. 5(a) and 5(e)], minimum of the second excitation gap $\Delta_{\text{ex}}^{(2)}$ [Figs. 5(b) and 5(f)] and sudden changes of structure factors [Figs. 5(c) and 5(g)] can be observed. Notice that in CDW phase the ground state and first excited state are almost degenerate. On the other hand, only “hump”s can be observed at the phase boundaries between the $C=2$ (or SDW) and the intermediate phase. Near the positions of the “hump”s, evolutions of the low-energy levels are more complicated and changes of structure factors are more smooth. Such features have also been discussed in Refs. Shao et al. (2021); Yuan et al. (2023), where its reason is attributed to the finite-size effect and it can be eliminated by using different clusters or twisted boundary conditions.

### v3

@@ III.1 Results of the exact diagonalization method
To elucidate the critical behaviors of the phase diagram in Fig. 1(a) with $U=0.0$, we focus on the two black dashed lines with $V=0.4$ and $\Delta=1.4$ to calculate the relevant properties. That is, the left and right panels in Fig. 2 are corresponding to the results of $V=0.4$ and $\Delta=1.4$, respectively. The first four lowest-lying energy levels $E_{\alpha}$ ($\alpha=0,1,2,3$ and $\alpha=0$ represents the ground state) are obtained by employing the Arnoldi Lehoucq et al. (1997) method, see Figs. 2(a) and 2(e). In Fig. 2(a), the system traverses the $C=2$, $C=1$ and SSO phases as $\Delta$ increases from $0.6$ to $1.4$, and two level crossings between the ground state and one excited state are observed at $\Delta\approx 0.9$ and $\Delta\approx 1.2$. In Fig. 2(e), the system crosses the SSO, $C=1$ and SCO phases as $V$ increases from $0.0$ to $2.0$, and two level crossings occur at $V\approx 0.75$ and $V\approx 1.25$. To provide a more intuitive representation, we exhibit the excitation gaps $\Delta_{\rm ex}^{(1)}=E_{1}-E_{0}$ and $\Delta_{\rm ex}^{(2)}=E_{2}-E_{0}$, as shown in Figs. 2(b) and 2(f). It is evident that the excitation gaps exhibit local minimum values at the critical points where level crossings occur. It is worth noting that in Fig. 2(f), $\Delta_{\rm ex}^{(1)}$ becomes very small when $V>1.25$, owing to the nearly degenerate ground state and first excited state in SCO phase. Another feature to characterize the phase transitions is the change of the SCO or SSO structure factor ($S_{\text{SCO}}$ or $S_{\text{SSO}}$), see Figs. 2(c) and 2(g). Values of $S_{\text{SCO}}$ ($S_{\text{SSO}}$) in $C=1$ phase is larger than those in $C=2$ phase but smaller than those in the SCO (SSO) phase, indicating again that $C=1$ is an intermediate state resulting from the interplay between topology, electronic correlations and lattice potentials. Finally, sharp peaks of the ground-state fidelity metrics $g$ in Figs. 2(d) and 2(h) can be observed and used to characterize the critical points. Besides, for $U=3.0$ in Fig. 1(d), similar calculations and discussions along the black dashed lines with $\Delta=0.4$ and $\Delta=0.6$ are detailed in Appendix B.
Before closing this section, we would like to note that the finite-size effect in ED method is discussed in Appendix C.
@@ III.2 Results of the mean-field method
To contrast the exact results on small lattice above, we now report the outcomes of the mean-field method. We use the $180\times 180$ lattice to calculate the band gap and the order parameters ($\mathcal{O}_{\text{SCO}}$ and $\mathcal{O}_{\text{SSO}}$), while a $20\times 20$ lattice is adopted for the calculations of Chern number. Despite the differences in lattice size, the critical points identified by the Chern number and other properties are consistent, as will be shown in Fig. 4. Phase diagrams with regard to the parameters $V$ and $\Delta$ are shown in Figs. 3(a), 3(b), 3(c) and 3(d) with $U=0.0$, $U=1.0$, $U=2.0$ and $U=3.0$, respectively. In comparison to the phase diagrams of ED in Fig. 1, similar features can be observed: SSO and SCO dominate the system for large $\Delta$ and $V$, respectively, leaving the $C=2$ phase in small ($V$, $\Delta$) region; the $C=1$ phase are surrounded by the above three phases. While differences between the MF and ED phase diagrams include:
(1) In the MF method, the increased interaction $U$ does not eliminate the $C=1$ phase at least for $U\leq 3.0$.
(2) The $C=1$ phase in Fig. 1(a) can exist with very small $V$, while in small $\Delta$ region ($\Delta\leq 0.5$) it vanishes. This is opposite to the MF results in Fig. 3(a), where the $C=1$ phase can exist in small $\Delta$ region rather than small $V$ region.
@@
Figure 3: Mean-field phase diagram in the parametric space ($V$, $\Delta$) of the model (1) based on the results of Chern number $C$ and the order parameters ($\mathcal{O}_{\text{SCO}}$ and $\mathcal{O}_{\text{SSO}}$), with (a) $U=0.0$, (b) $U=1.0$, (c) $U=2.0$ and (d) $U=3.0$.
@@
Taking the $U=0.0$ case in Fig. 3(a) as an example, we show in Figs. 4(a), 4(b), 4(c) and 4(d) the results of the SCO order parameter ($\mathcal{O}_{\text{SCO}}$), the SSO order parameter ($\mathcal{O}_{\text{SSO}}$), the band gap ($\Delta(\textbf{k})$) and the Chern number ($C$), respectively, as a function of $V$ and $\Delta$. It is noteworthy that the phase boundary between SSO and SCO can only be identified from the order parameters, and other phase boundaries are characterized by the results of Chern number, as seen in Fig. 4(d). Despite some defect points in Fig. 4(d), possibly attributed to the severe quantum fluctuation near the critical points, the phase boundaries can be readily identified. In Fig. 4(a), we observe that $\mathcal{O}_{\text{SCO}}$ vanishes in the SSO and $C=2$ phases, while the $C=1$ phase can be regarded as a zone of transition with $\mathcal{O}_{\text{SCO}}$ changing from $0.0$ to $2.0$. For $\mathcal{O}_{\text{SSO}}$ in Fig. 4(b), the $C=1$ phase remains a transition region with finite values but smaller that those in SSO phase. These features are similar to the structure factors of ED results shown in Figs. 2(c) and 2(g). To contrast the excitation gaps in ED results, we show the band gap defined as $\Delta(\textbf{k})$=min [$E_{2}(\textbf{k}^{\prime})-E_{1}(\textbf{k}^{\prime})$] in Fig. 4(c). The rapid drop or closing of the gap size can be observed at the critical values, except for the phase boundary between SSO and SCO, where only discontinuity in gap size appears. This can be connected with the general picture that the change of a topological invariant is always accompanied by a single-particle gap closing. While for the phase transition from SSO to SCO side, the Chern number does not change ($C=0$). This may explain the discontinuity of gap size instead of gap closing, considering that the SSO and SCO states are all gapped phases with different gap sizes.
@@
Figure 4: Contour plots of (a) the SCO order parameter, (b) the SSO order parameter, (c) the band gap and (d) the Chern number, as a function of $V$ and $\Delta$. The mean-field approach is adopted for the calculations of the model (1) with $U=0.0$.
To analyze the origin of the $C=1$ phase, we define the effective potential differences $\Delta_{\rm MF}^{\uparrow}$ and $\Delta_{\rm MF}^{\downarrow}$ in Eq. (13), and here we show $\Delta_{\rm MF}^{\uparrow}/t_{2}$ and $\Delta_{\rm MF}^{\downarrow}/t_{2}$ as a function of $V$ in Fig. 5(a). We set $U=0$, $\Delta=0.6$ and $V\in[0.0,1.5]$, see the black dashed line in Fig. 3(a). It is known that in the noninteracting spinless Haldane model, the Chern number $C$ equals $1$ under the condition $|\Delta|/t_{2}<3\sqrt{3}$, and $C$ equals 00 when $|\Delta|/t_{2}>3\sqrt{3}$. From Fig. 5(a) we can observe that:
$\Delta_{\rm MF}^{\uparrow}/t_{2}<3\sqrt{3}$ and $\Delta_{\rm MF}^{\downarrow}/t_{2}<3\sqrt{3}$ when $C=2$;
$\Delta_{\rm MF}^{\uparrow}/t_{2}>3\sqrt{3}$ and $\Delta_{\rm MF}^{\downarrow}/t_{2}<3\sqrt{3}$ when $C=1$;
$\Delta_{\rm MF}^{\uparrow}/t_{2}>3\sqrt{3}$ and $\Delta_{\rm MF}^{\downarrow}/t_{2}>3\sqrt{3}$ when $C=0$.
Here the vertical dashed lines in Fig. 5(a) separate the $C=2$, $C=1$ and $C=0$ regions. In Fig. 5(b), we show the band gap $\Delta(\textbf{k})$ as a function of $V$, where two gap closures can be found at the critical points, denoted by the vertical dashed lines identical to those in Fig. 5(a). These features indicates that the formation of the $C=1$ phase is due to the spontaneous SU(2) symmetry breaking, with only one spin species in topological state.
Figure 5: (a) The effective potential differences in units of $t_{2}$ $\Delta_{\rm MF}^{\uparrow}/t_{2}$ and $\Delta_{\rm MF}^{\downarrow}/t_{2}$, and (b) the band gap as a function of $V\in[0.0,1.5]$ (see the black dashed line in Fig. 3(a)). Other parameters: $U=0$ and $\Delta=0.6$. The horizontal dashed line in (a) denotes the value of $3\sqrt{3}$ in y axis, and the vertical dashed lines in (a) and (b) separate the $C=2$, $C=1$ and $C=0$ regions.
@@ IV Summary and Discussion
In summary, we investigated the interacting spinful Haldane model at half-filling on the honeycomb lattice with spin-dependent sublattice potentials. By employing the exact-diagonalization (ED) and mean-field (MF) methods, we obtained similar ground-state phase diagrams for the nearest-neighbor interaction $V\in[0,3]$ and sublattice potential difference $\Delta\in[0,3]$. The staggered spin order (SSO) and staggered charge order (SCO) phases dominate in the large $\Delta$ and $V$ region, respectively; the $C=2$ phase prevails when both $V$ and $\Delta$ are small enough; and an intermediate phase with $C=1$ is surrounded by the aforementioned three phases. Except for the change in the topological invariant, other features such as the closure of the excitation gap, changes in structure factors, and peaks in the fidelity metric were also observed in the ED results. Meanwhile, we examined the mean-field Chern number, band gap and local order parameters, also providing clear evidence for the existence of the $C=1$ phase. Especially, by analyzing the effective potential differences $\Delta_{\rm MF}^{\uparrow}$ and $\Delta_{\rm MF}^{\downarrow}$, the origin of this exotic phase is also attributed to the spontaneous SU(2) symmetry breaking.
However, the interaction $U$ suppresses the $C=1$ phase, raising an open question about its potential existence as $U$ increased to $3.0$. In the ED approach, $C=1$ can not be found when $U=3.0$ even though the “intermediate phase” can still be identified from the fidelity metric results. In the MF approach, the $C=1$ phase preserves for $U=3.0$, albeit with a very small area.
Finally, we would like to note that in previous investigations, the staggered spin order (SSO) has been denoted as the antiferromagnetic (AF)Jiang et al. (2018); Wang and Qi (2019) or spin-density-wave (SDW) state He et al. (2011, 2024); Silva et al. ; Yuan et al. (2023); Shao et al. (2021), and the staggered charge order (SCO) has been labeled as the band insulator (BI)Wang and Qi (2019); Vanhala et al. (2016) or charge-density-wave (CDW)Wang and Qi (2019); He et al. (2024); Silva et al. ; Yuan et al. (2023); Shao et al. (2021) state. While in this paper, we refer to them as SSO and SCO to stress their staggered characteristics on A and B sublattices.
@@ Acknowledgements.
The authors acknowledge insightful discussions with E. V. Castro, R. Mondaini, H. Lu and S. Hu. C. S. acknowledges support from the National Natural Science Foundation of China (NSFC; Grant No. 12104229) and the Fundamental Research Funds for the Central Universities (Grant No. 30922010803). H.-G. L. acknowledges support from NSFC (Grants No. 11834005 and No. 12247101), and the National Key Research and Development Program of China (Grant No. 2022YFA1402704).
Appendix A the ED phase diagram in the parametric space ($U$, $\Delta$) with $V=0$
Figure 6: (a) The exact-diagonalization (ED) phase diagram of the model (1) in the parametric space($U$, $\Delta$) with $V=0$. The black dashed line denotes the parameters we choose to show the structure factor $S_{\text{SSO}}$ in (b), the excitation gaps $\Delta_{\text{ex}}^{(\alpha)}$ in (c), and the fidelity metric $g$ in (d).
To better understand the impact of the interaction parameter $U$ on the phase diagram, we present the ED phase diagram of our model as a function of $U$ and $\Delta$ in Fig. 6(a), with $V=0$. It can be observed that both $U$ and $\Delta$ favor the SSO order, and $\Delta$ is more efficient in suppressing the $C=2$ phase compared to $U$. Specifically, for $\Delta=0$, the critical point is located at $U=8.1$, while for $U=0$, the critical point is located at $\Delta=1.03$. This explains why increasing $U$ from 00 to $3$ does not eliminate the intermediate phase in the ($V$, $\Delta$) phase diagram, as depicted in Fig. 1 of the main text.
Furthermore, we choose $U=4$ [black dashed line in Fig. 6(a)] to plot the structure factor of SSO in Fig. 6(b), the excitation gaps in Fig. 6(c) and the fidelity metric in Fig. 6(d). The rapid change of $S_{\text{SSO}}$, the closure of $\Delta_{\text{ex}}^{(1)}$ and the sharp peak of the fidelity metric $g$ can be observed, respectively. Notably, no intermediate phase can be identified from these properties, indicating that the $C=1$ phase results from the competition between $V$ and $\Delta$.
Appendix B details of some properties in ED method with $U=3$
@@
Figure 7: (a)(e) Four lowest-lying energy levels $E_{\alpha}$, (b)(f) the excitation gaps $\Delta_{\text{ex}}^{(\alpha)}$, (c)(g) the structure factors $S_{\text{SSO/SCO}}$, and (d)(h) the fidelity metric $g$ of the model (1) with $\Delta=0.4$ on the left panels and $\Delta=0.6$ on the right panels. The on-site interaction $U=3$ and the parameters are corresponding to the black dashed lines in Fig. 1(d).
@@
In Fig. 7, considering the case of $U=3.0$, we choose $\Delta=0.4$ (left panel) and $\Delta=0.6$ (right panel) to illustrate the relevant properties. The line of $\Delta=0.4$ ($\Delta=0.6$) in Fig. 1(d) crosses the $C=2$ phase (SSO phase), the “intermediate phase” and the SCO phase as $V$ increases from $1.0$ to $2.0$ [see black dashed lines in Fig. 1(d)]. Examining the fidelity metrics $g$ in Figs. 7(d) and 7(h), we observe sharp peaks between the “intermediate phase” and the SCO phase, located at $V\approx 1.5$ and $V\approx 1.6$, respectively. In the vicinity of the locations of the shape peaks, level crossings between the ground state and the second excited state can be observed [Figs. 7(a) and 7(e)], as well as the minimum values of the second excitation gap $\Delta_{\text{ex}}^{(2)}$ [Figs. 7(b) and 7(f)] and sudden changes in structure factors [Figs. 7(c) and 7(g)]. Notice that in SCO phase the ground state and first excited state are nearly degenerate.
On the other hand, only “hump”s are observed at the phase boundaries between the $C=2$ (or SSO) and the “intermediate phase”, locating at $V\approx 1.32$ in Fig. 7(d) [$V\approx 1.27$ in Fig. 7(h)]. Near the positions of the “hump”s, level crossings between the ground state and the first excited state [Figs. 7(a) and 7(e)], the local minimum values of the first excitation gap $\Delta_{\text{ex}}^{(1)}$ [Figs. 7(b) and 7(f)] and smooth changes in structure factors [Figs. 7(c) and 7(g)] can be observed. Similar features have been discussed in Refs. Shao et al. (2021); Yuan et al. (2023), where the reason is attributed to the finite-size effect, and it can be mitigated by employing different clusters or twisted boundary conditions.
Appendix C the ED phase diagram in the parametric space ($V$, $\Delta$) on a 6-site lattice
Figure 8: (a) The exact-diagonalization (ED) phase diagram of the model (1) in the parametric space($V$, $\Delta$) with $U=0$, on a $6$-site cluster shown in (c). The black dashed line in (a) denotes the parameters we choose to show the spectrum in (b), the excitation gap $\Delta_{\text{ex}}^{(2)}$ in (c), and the structure factors $S_{\rm SSO/SCO}$ in (d).
In the beginning of Sec. III, we mentioned that we utilized a $12$A cluster, whose reciprocal lattice contains the K points Shao et al. (2021), to study the phase transitions of the model (1) in the main text. Now we would like to discuss the finite-size effect in our ED study. Is is known that during the phase transition of noninteracting Haldane model, the gap closes at the K points. It’s important to note that the condition for reciprocal lattice of a cluster to contain the K points can only be satisfied when the site number of the cluster is a multiple of $6$Varney et al. (2010). This condition has been demonstrated to be crucial for characterizing the topological phase transitions in the interacting Haldane modelsVarney et al. (2010, 2011); Shao et al. (2021); Yuan et al. (2023). Unfortunately, due to limitations in our current computational resources, we are unable to access an $18$A cluster. Therefore, we conducted our phase diagram analysis on the $6$A cluster instead, as shown in subplot of Fig. 8(c).
As expected, an intermediate phase can be observed in the ($V$, $\Delta$) phase diagram in Fig. 8(a) and we select the dashed line with $\Delta=1.5$ to illustrate more detailed properties. In Fig. 8(b), we depict four lowest-lying energy levels, and the level crossings at two critical points can be observed. It’s worth noting that the ground state $E_{0}$ and first excited state $E_{1}$ are degenerate in the SCO and intermediate phase, while the first excited state $E_{1}$ and the second excited state $E_{2}$ are degenerate in the SSO phase. As a result, it is more appropriate to use the second excitation gap, $\Delta_{\rm ex}^{(2)}$, to characterize the gap closure at phase transition points, as shown in Fig. 8(c). Additionally, we present the structure factors $S_{\rm SSO}$ and $S_{\rm SCO}$ in Fig. 8(d), where the transition nature of the intermediate phase is clearly visible.